# Charged current neutrino scattering from nucleons

### v1

@@ I Introduction
  • Neutrino Mass Hierarchy: Determining the precise order of neutrino masses—whether the mass states follow a normal hierarchy ($m_{1}~{}<~{}m_{2}~{}<~{}m_{3}$) or an inverted hierarchy ($m_{3}~{}<~{}m_{1}~{}<~{}m_{2}$) remains a critical unanswered question. Upcoming high-precision experiments like the Deep Underground Neutrino Experiment (DUNE) at the Fermilab in the United States and Hyper-Kamiokande in Japan aim to resolve this by studying differences in the neutrino oscillation patterns.
  • Absolute Neutrino Mass Scale: Although we know neutrinos have mass from the oscillation experiments, the absolute scale of these masses is still unknown. The KATRIN experiment seeks to measure the electron neutrino’s mass directly through the beta decay of tritium. However, the tiny mass of neutrinos makes this measurement extremely challenging, and determining the masses of other neutrino flavors is even more difficult as currently the measurements are being done in the pion and tauon decay experiments.
@@
  • CP Violation in the Lepton Sector: While CP violation (which breaks the symmetry between matter and antimatter) has been observed in the hadron sector, it has not yet been confirmed in the lepton sector. Recent studies have placed limits on the CP-violating phase $\delta$, but more precise experiments are needed. Discovering CP violation in neutrino oscillations could provide insights into why the universe is dominated by matter. Experiments like DUNE and Hyper-Kamiokande are focused on measuring CP-violating phases in the neutrino oscillations.
  • Sterile Neutrinos: The LSND experiment at the Los Alamos using $\mu$-decay at rest and observing $\bar{\nu}_{e}$ appearance in $\bar{\nu}_{\mu}$, followed by the MiniBooNE accelerator experiment using $\nu_{e}$/$\bar{\nu}_{e}$ appearance in $\nu_{\mu}$/$\bar{\nu}_{\mu}$ beam at the Fermilab have observed excess of electron events which can not be explained in the standard three flavor neutrino oscillation scenario and hints towards a possible fourth flavor (sterile) neutrino, which does not interact with the matter in the ordinary way like the other three neutrinos. The recent experiments like the MicroBooNE, IceCube, STEREO, etc. have refuted the idea of sterile neutrino. Further investigations are necessary to confirm or deny the presence of sterile neutrinos.
@@
  • High-Energy Neutrino Astronomy: The detection of high-energy neutrinos from the astrophysical sources by the experiments like the IceCube, has opened a new window into the universe. There are other experiments like PINGU and KM3NeT to observe the high energy neutrinos. However, understanding the sources and mechanisms producing these neutrinos remains a significant challenge, both theoretically and experimentally.
@@
With the theoretical prediction Lee:1956vjd ; Lee:1956qn and the experimental observation of parity violation Wu:1957my in $\beta$ decays, the experimental determination of the neutrino helicity to be $-1$ Goldhaber:1958nb , and the longitudinal polarization of $e^{-}~{}(e^{+})$ in $\beta^{-}~{}(\beta^{+})$ decays to be $\mp\beta$ Frauenfelder:1957na ; Page:1957zza , the $V-A$ theory of weak interactions was proposed by Sudarshan and Marshak Sudarshan:1958vf , Feynman and Gell-Mann Feynman:1958ty , and Sakurai Sakurai:1958zz , using chiral($\gamma_{5}$) invariance Salam:1957st ; Landau:1957tp ; Lee:1957qr , in which the interaction Lagrangian is given by
@@
In the standard model, the interaction Lagrangian for the weak charged current (CC) and neutral current (NC) are written as:
@@
In analogy with the Lagrangian for the weak interaction for the leptons, the most general Lagrangian for the weak charged current interaction, for the 4-quark flavor, is written as
@@
Similarly, the neutral current weak interaction Lagrangian for the quarks is written as:
@@
States $g_{V}$ $g_{A}$
$u$ $\frac{1}{2}-\frac{4}{3}\sin^{2}\theta_{W}$  $1/2$
$d$ $-\frac{1}{2}+\frac{2}{3}\sin^{2}\theta_{W}$ $-1/2$
Table 1: Couplings of the quarks($u,d$) to $Z_{\mu}$ field.
However, certain phenomena, such as the neutrino mixing and oscillations–which imply that neutrinos possess small masses—as well as observed decay processes involving flavor-changing neutral currents (FCNC) and lepton flavor violation (LFV), remain unexplained by the Standard Model and suggest the need for the physics Beyond the Standard Model (BSM).
In this paper, we have studied the charged current induced (anti)neutrino scattering from the free nucleon target. In the (anti)neutrino energy region of a few GeV, relevant for most of the present and future neutrino oscillation experiments using the accelerator neutrinos like T2K, MicroBooNE, NOvA, DUNE, T2HK, etc., the neutrino cross section receives contribution from the quasielastic, inelastic and deep inelastic scattering processes. In Section II, we discuss in brief the formalism of the aforementioned processes. In the case of inelastic scattering processes, we have studied the production of strange baryon like $\Lambda$, $\Sigma$, and mesons like pion, kaon, $\eta$ meson, in the final state as well as the associated particle production. The results for the $Q^{2}$ distribution of the cross section for these processes and their discussions are presented in Section III. Section IV concludes the findings of the present work.
@@ II.1 Quasielastic scattering
In the above expression, $Q^{2}=-q^{2}=-(k-k^{\prime})^{2}$ is the four momentum transfer squared, $M$ and $M^{\prime}$ are the masses of the initial and the final baryons, respectively. $f_{1}^{NN^{\prime}(Y)}(Q^{2})$, $f_{2}^{NN^{\prime}(Y)}(Q^{2})$ and $f_{3}^{NN^{\prime}(Y)}(Q^{2})$ are the vector, weak magnetic and induced scalar $N-N^{\prime}$ or $N-Y$ transition form factors corresponding to the $\Delta S=0$ or $|\Delta S|=1$ processes, respectively, and $g_{1}^{NN^{\prime}(Y)}(Q^{2})$, $g_{2}^{NN^{\prime}(Y)}(Q^{2})$ and $g_{3}^{NN^{\prime}(Y)}(Q^{2})$ are the axial-vector, induced tensor (or weak electric) and induced pseudoscalar form factors, respectively. These form factors have been determined using the various symmetry properties of the weak hadronic currents like the T invariance, G invariance, conserved vector current hypothesis, etc., and the details can be found in Refs. SajjadAthar:2022pjt ; Fatima:2018tzs .
@@ II.2 Inelastic scattering
where in Eq. (23), $m(=\pi,\eta,K$, etc.) is a meson produced with a baryon ($B=N,Y$, etc.) in the final state. $k(k^{\prime})$ is the four momentum of the incoming (outgoing) lepton having energy $E_{\nu}(E_{l})$; $p$ is the four momentum of the incoming nucleon which is at rest, $E_{B}$ and $p^{\prime}$ are respectively the energy and four momentum of the outgoing baryon, and the meson four momentum is $p_{m}$ with energy $E_{m}$, and $M$ is the nucleon mass. ¯$\overline{\sum}\sum|\mathcal{M}|^{2}$ is the square of the transition amplitude averaged (summed) over the spins of the initial (final) states and the transition matrix element is written in terms of the leptonic and the hadronic currents as
@@
[…]
where $W~{}(=(p+q)^{2})$ is the center-of-mass (CM) energy, $E_{m}^{min}$ and $E_{m}^{max}$ are respectively the minimum and maximum energy of the outgoing meson and $\phi_{qp_{m}}$ is the azimuthal angle between the scattering plane and the reaction plane as shown in Fig.1.
Figure 2: Feynman diagram for CC inelastic scattering processes.
@@ II.3 Deep inelastic scattering (DIS)
$\frac{d^{2}\sigma}{dxdy}=\frac{yM}{\pi}~{}\frac{E_{\nu}}{E_{l}}~{}\frac{|{\vec{k}^{\prime}}|}{|{\vec{k}}|}\;\overline{\sum}\sum|{\cal M}|^{2}.$ (29)
@@
For the present numerical calculations, we have used the MMHT nucleonic PDFs parameterization Harland-Lang:2014zoa in the MSbar scheme.
@@ III.1 Quasielastic scattering
In Figs. 4 and 5, we have presented the results for the $Q^{2}$-distribution at $E_{{\nu_{\mu}}({\bar{\nu}_{\mu}})}=2$ GeV i.e. $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced quasielastic scattering processes on the neutron ($\nu_{\mu}+n\rightarrow\mu^{-}+p$) and proton ($\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+n$) targets, respectively. These results are presented using the BBBA05 Bradford:2006yz parameterization of the Sachs’ electric and magnetic form factors for the isovector vector form factors $f_{1,2}(Q^{2})$, axial vector form factor having dipole parameterization i.e. $g_{1}(Q^{2})=\frac{g_{1}(0)}{\left(1+\frac{Q^{2}}{M_{A}^{2}}\right)^{2}}$, $g_{1}(0)~{}(=1.26)$ is the axial charge and $M_{A}$ is the axial dipole mass taken to be the world average value i.e. $M_{A}=1.026$ GeV. In the charged current sector, where the partial conservation of axial vector current (PCAC) is assumed to be valid, the pseudoscalar form factor $g_{3}(Q^{2})$, is dominated by the pion pole dominance of the divergence of the axial vector current (PDDAC) and is given in terms of $g_{1}(Q^{2})$ using the Goldberger-Trieman relation. However, $g_{3}(Q^{2})$ does not contribute in the limit $m_{l}\rightarrow 0$. The form of the induced tensor form factor $g_{2}(Q^{2})$, which is associated with the second class currents, is also taken to be of dipole form with $g_{2}(0)=\pm 3$ and same $M_{A}$ as for the $g_{1}(Q^{2})$. The results presented in Fig. 4 are obtained to study the effect of second class currents assuming T-invariance by varying $g_{2}^{R}(0)$ between 00 and ±$\pm 3$, while the results have been presented in Fig. 5 to study the effect of second class currents assuming time reversal violation by varying $g_{2}^{I}(0)$ between 00 and ±$\pm 3$.
It may be observed from the figure that in the presence of second class current(SCC) assuming time reversal invariance (Fig. 4), which requires $g_{2}(Q^{2})$ to be purely real i.e. $g_{2}(Q^{2})=g_{2}^{R}(Q^{2})$, the cross sections increases with the increase in the value of $g_{2}^{R}(0)$ and a change in the value of $g_{2}^{R}(0)$ from 0 to 3 results in an increase in the cross section by $25\%$ at $Q^{2}\sim 0.4-0.6$ GeV2, for both neutrino and antineutrino induced reactions. We find that a change of sign in $g_{2}(0)$ i.e. $g_{2}(0)=-3$ gives almost the same result $g_{2}(0)=+3$. In the case of second class currents assuming time reversal violation, which requires $g_{2}(Q^{2})$ to be purely imaginary i.e. $g_{2}(Q^{2})=g_{2}^{I}(Q^{2})$, a similar effect on the cross section (Fig. 5) is observed, as seen in the case of second class currents assuming time reversal invariance.
@@
In Figs. 6 and  7, we have presented the results for the $Q^{2}$-distribution at $E_{\bar{\nu}_{\mu}}=2$ GeV for the charged current $\bar{\nu}_{\mu}$ induced $|\Delta S|=1$ quasielastic scattering processes on the proton ($\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+\Lambda$) (left panel) and neutron ($\bar{\nu}_{\mu}+n\rightarrow\mu^{+}+\Sigma^{-}$) (right panel) targets, respectively. Assuming SU(3) symmetry, the $N-Y(=\Lambda,\Sigma)$ transition form factors have been taken to be of the similar form as taken in the case of the $|\Delta S|=0$ quasielastic scattering processes and discussed above. The results presented in Fig. 6 are obtained in the presence of the second class currents assuming time reversal invariance by varying $g_{2}^{R}(0)$ between 00 and ±$\pm 3$, while the results of T-violation studies have been presented in Fig. 7 by varying $g_{2}^{I}(0)$ between 00 and ±$\pm 3$. It may be observed from the figures that the non-zero value of the form factor associated with the second class currents assuming time reversal invariance as well as its violation, increases the cross section, which is more pronounced in the case of $\Lambda$ production as compared to the results obtained for $\Sigma^{-}$ production. Unlike the case of $\Delta S=0$ quasielastic processes, in the case of $\Lambda$ production, we find that the cross section obtained using the negative value of $g_{2}^{R}(0)$ is larger than the one obtained using the positive value of $g_{2}^{R}(0)$. Quantitatively, in the case of $\Lambda$ production, we find an enhancement of about $25\%$ ($40\%$) at $Q^{2}=0.5$ GeV2 when $g_{2}^{R}(0)=+3~{}(-3)$ is taken in the numerical calculations, and this enhancement increases with the increase in the value of $Q^{2}$. It may be pointed out that the $\Lambda$ production is more suitable for the SCC or T-violation studies as their effect on $\Lambda$ polarization can be studied by analyzing the angular distribution of pions produced in $\Lambda\rightarrow p\pi^{-}$ decaysSajjadAthar:2022pjt .

### v2

@@ I Introduction
  • Neutrino Mass Hierarchy: Determining the precise order of neutrino masses—whether the mass states follow a normal hierarchy ($m_{1}~{}<~{}m_{2}~{}<~{}m_{3}$) or an inverted hierarchy ($m_{3}~{}<~{}m_{1}~{}<~{}m_{2}$) remains a critical unanswered question. Upcoming high-precision experiments like the Deep Underground Neutrino Experiment (DUNE) DUNE:2018tke ; DUNE:2024wvj at the Fermilab in the United States, Hyper-Kamiokande Hyper-Kamiokande:2018ofw in Japan, The Jiangmen Underground Neutrino Observatory (JUNO) Cerrone:2024vvk in China, etc., aim to resolve this by studying differences in the neutrino oscillation patterns.
  • Absolute Neutrino Mass Scale: Although we know neutrinos have mass from the oscillation experiments, the absolute scale of these masses is still unknown. The KATRIN experiment KATRIN:2022ayy seeks to measure the electron neutrino’s mass directly through the beta decay of tritium. However, the tiny mass of neutrinos makes this measurement extremely challenging, and determining the masses of other neutrino flavors is even more difficult as currently the measurements are being done in the pion and tauon decay experiments.
@@
  • CP Violation in the Lepton Sector: While CP violation (which breaks the symmetry between matter and antimatter) has been observed in the hadron sector, it has not yet been confirmed in the lepton sector. Recent studies have placed limits on the CP-violating phase $\delta$, but more precise experiments are needed. Discovering CP violation in neutrino oscillations could provide insights into why the universe is dominated by matter. Experiments like DUNE DUNE:2018tke ; DUNE:2024wvj and Hyper-Kamiokande Hyper-Kamiokande:2018ofw are focused on measuring CP-violating phases in the neutrino oscillations.
  • Sterile Neutrinos: The LSND experiment LSND:1996ubh ; LSND:2001aii at the Los Alamos using $\mu$-decay at rest and observing $\bar{\nu}_{e}$ appearance in $\bar{\nu}_{\mu}$, followed by the MiniBooNE MiniBooNE:2007uho ; MiniBooNE:2008yuf accelerator experiment using $\nu_{e}$/$\bar{\nu}_{e}$ appearance in $\nu_{\mu}$/$\bar{\nu}_{\mu}$ beam at the Fermilab have observed excess of electron events which can not be explained in the standard three flavor neutrino oscillation scenario and hints towards a possible fourth flavor (sterile) neutrino, which does not interact with the matter in the ordinary way like the other three neutrinos. The recent experiments like the MicroBooNE Denton:2021czb , IceCube IceCube:2024dlz , STEREO STEREO:2022nzk , etc. have refuted the idea of sterile neutrino. Further investigations are necessary to confirm or deny the presence of sterile neutrinos.
@@
  • High-Energy Neutrino Astronomy: The detection of high-energy neutrinos from the astrophysical sources by the experiments like the IceCube IceCube:2024dlz , has opened a new window into the universe. There are other experiments like PINGU IceCube:2016xxt and KM3NeT Ferrara:2024try to observe the high energy neutrinos. However, understanding the sources and mechanisms producing these neutrinos remains a significant challenge, both theoretically and experimentally.
@@
With the theoretical prediction Lee:1956vjd ; Lee:1956qn and the experimental observation of parity violation Wu:1957my in $\beta$ decays, the experimental determination of the neutrino helicity to be $-1$ Goldhaber:1958nb , and the longitudinal polarization of $e^{-}~{}(e^{+})$ in $\beta^{-}~{}(\beta^{+})$ decays to be $\mp\beta$ Frauenfelder:1957na ; Page:1957zza , the $V-A$ theory of weak interactions was proposed by Sudarshan and Marshak Sudarshan:1958vf , Feynman and Gell-Mann Feynman:1958ty , and Sakurai Sakurai:1958zz , using chiral ($\gamma_{5}$) invariance Salam:1957st ; Landau:1957tp ; Lee:1957qr , in which the interaction Lagrangian is given by Athar:2020kqn :
@@
In the standard model, the interaction Lagrangian for the weak charged current (CC) and neutral current (NC) are written as Athar:2020kqn :
@@
States $g_{V}$ $g_{A}$
$u$ $\frac{1}{2}-\frac{4}{3}\sin^{2}\theta_{W}$  $1/2$
$d$ $-\frac{1}{2}+\frac{2}{3}\sin^{2}\theta_{W}$ $-1/2$
Table 1: Couplings of the quarks ($u,d$) to $Z_{\mu}$ field.
In analogy with the Lagrangian for the weak interaction for the leptons, the most general Lagrangian for the weak charged current interaction, for the 4-quark flavor, is written as Athar:2020kqn :
@@
Figure 1: Kinematical $Q^{2}-W$ plane depicting neutrino-nucleon scattering at two representative laboratory neutrino energies, i.e. 2 GeV (upper panel) and 3 GeV (lower panel).
Similarly, the neutral current weak interaction Lagrangian for the quarks is written as Athar:2020kqn :
@@
However, certain phenomena, such as the neutrino mixing and oscillations–which imply that neutrinos possess small masses–as well as observed decay processes involving flavor-changing neutral currents (FCNC) and lepton flavor violation (LFV), remain unexplained by the Standard Model and suggest the need for the physics Beyond the Standard Model (BSM).
In this paper, we have studied the charged current induced (anti)neutrino scattering from the free nucleon target. In the (anti)neutrino energy region of a few GeV, relevant for most of the present and future neutrino oscillation experiments using the accelerator neutrinos like T2K, MicroBooNE, NOvA, DUNE, T2HK, etc., the neutrino cross section receives contribution from the quasielastic, inelastic and deep inelastic scattering processes. In general, when an (anti)neutrino interacts with a nucleon target, then the center of mass energy $W$, given by $W^{2}=(p+q)^{2}$, where $q$ is the four-momentum transferred to the hadronic system, corresponds to the quasielastic scattering when $W=M$. As $W$ increases, the processes involving the single pion, multiple pion, kaon, eta and other meson production starts contributing and finally to the deep inelastic scattering (DIS) region, where jet of hadrons are produced in the final state. The intermediate region, spanning the meson production threshold to the DIS domain, is referred to in the literature as the shallow inelastic scattering (SIS) region. The kinematic region defining the SIS and the onset of DIS is not free from ambiguities, and to define these two kinematic regions the experiments use different cuts on $W$. To kinematically depict the different scattering processes relevant for the neutrino-nucleon cross sections in the few GeV energy region, in Fig. 1, we have shown the $Q^{2}~{}(=-q^{2}\geq 0)$ vs. $W$ plot at two different values of neutrino energies viz. 2 and 3 GeV. It may be observed from the figure that with the increase in $W$, several resonances contribute and besides pions, many other inelastic processes also start contributing to the neutrino-nucleon cross sections.
In Section II, we discuss in brief the formalism of the aforementioned processes. In the case of inelastic scattering processes, we have studied the production of strange baryon like $\Lambda$, $\Sigma$, and mesons like pion, kaon, $\eta$ meson, in the final state as well as the associated particle production in which kaons and hyperons both are produced. The results for the $Q^{2}$ distribution of the cross section for these processes and their discussions are presented in Section III. Section IV concludes the findings of the present work.
@@ II.1 Quasielastic scattering
In the above expression, $Q^{2}=-q^{2}=-(k-k^{\prime})^{2}$ is the four momentum transfer squared, $M$ and $M^{\prime}$ are the masses of the initial and the final baryons, respectively. $f_{1}^{NN^{\prime}(Y)}(Q^{2})$, $f_{2}^{NN^{\prime}(Y)}(Q^{2})$ and $f_{3}^{NN^{\prime}(Y)}(Q^{2})$ are the vector, induced weak magnetic and induced scalar $N-N^{\prime}$ or $N-Y$ transition form factors corresponding to the $\Delta S=0$ or $|\Delta S|=1$ processes, respectively, and $g_{1}^{NN^{\prime}(Y)}(Q^{2})$, $g_{2}^{NN^{\prime}(Y)}(Q^{2})$ and $g_{3}^{NN^{\prime}(Y)}(Q^{2})$ are the axial-vector, induced tensor (or weak electric) and induced pseudoscalar form factors, respectively. These form factors have been determined using the various symmetry properties of the weak hadronic currents like the time-reversal (T) invariance, G-parity (G) invariance, conserved axial-vector current (CVC) and partially conserved axial-vector current (PCAC) hypotheses, etc., and the details can be found in Refs. SajjadAthar:2022pjt ; Fatima:2018tzs . The time reversal invariance implies that all the vector and axial vector form factors are real. G-parity is a multiplicative quantum number that results from the generalization of C-parity to multiplets of mesons (like pions, rho, etc.), and G-parity${}(G=Ce^{i\pi I_{y}})$ is defined as a combination of charge conjugation (C) and a 180 rotation around the y-axis of isospin space ($I_{y})$. The concept of G-invariance was first introduced by Weinberg to classify the above mentioned form factors as first and second class current form factors. The first class vector and axial vector form factors are those form factors that transform in the same way as the basic form factors $f_{1}(Q^{2})$ and $g_{1}(Q^{2})$ do, while the second class current form factors are those that transform with an opposite sign in their respective currents. With this definition the form factors $f_{1}(Q^{2})$, $f_{2}(Q^{2})$, $g_{1}(Q^{2})$, and $g_{3}(Q^{2})$ are classified as first class currents while $f_{3}(Q^{2})$ and $g_{2}(Q^{2})$ are classified as the second class current form factors. If G invariance is assumed to be an exact symmetry, then the form factors associated with the second class currents vanish. In the vector sector, the form factor $f_{3}(Q^{2})$ is further constrained to vanish due to CVC. Therefore, to study the effect of G-invariance one takes $g_{2}(Q^{2})$ to be non-zero in the numerical calculations. Moreover, if one assumes $g_{2}(Q^{2})$ to be purely real, then it preserves T invariance, while the complex or purely imaginary values of $g_{2}(Q^{2})$ results in T violation in the weak sector along with G violation.
@@ II.2 Inelastic scattering
where in Eq. (23), $m~{}(=\pi,\eta,K$, etc.) is a meson produced with a baryon ($B=N,Y$, etc.) in the final state. $k(k^{\prime})$ is the four momentum of the incoming (outgoing) lepton having energy $E_{\nu}(E_{l})$; $p$ is the four momentum of the incoming nucleon which is at rest, $E_{B}$ and $p^{\prime}$ are respectively the energy and four momentum of the outgoing baryon, and the meson four momentum is $p_{m}$ with energy $E_{m}$, and $M$ is the nucleon mass. ¯$\overline{\sum}\sum|\mathcal{M}|^{2}$ is the square of the transition amplitude averaged (summed) over the spins of the initial (final) states and the transition matrix element is written in terms of the leptonic and the hadronic currents as
@@
[…]
where $E_{m}^{min}$ and $E_{m}^{max}$ are respectively the minimum and maximum energy of the outgoing meson and $\phi_{qp_{m}}$ is the azimuthal angle between the scattering plane and the reaction plane as shown in Fig.2.
Figure 3: Feynman diagram for CC inelastic scattering processes.
@@ II.3 Deep inelastic scattering (DIS)
$\frac{d^{2}\sigma}{dxdy}=\frac{yM}{\pi}~{}\frac{E_{\nu}}{E_{l}}~{}\frac{|{\vec{k}^{\prime}}|}{|{\vec{k}}|}\;\overline{\sum}\sum|{\cal M}|^{2},$ (29)
where the four momentum transfer $q=k-k^{\prime}$, the Bjorken variable $x=\frac{Q^{2}}{2p\cdot q}$ and the inelasticity $y=\frac{p\cdot q}{p\cdot k}$ with $Q^{2}(\geq 0)=-q^{2}$.
@@
Figure 5: $\frac{d\sigma}{dQ^{2}}$ vs. $Q^{2}$ for the processes $\nu_{\mu}+n\longrightarrow\mu^{-}+p$ (left panel) and $\bar{\nu}_{\mu}+p\longrightarrow\mu^{+}+n$ (right panel) at $E_{\nu}=2$ GeV at different values of $M_{A}$ viz. $M_{A}=1.026$ GeV (solid line), 1.1 GeV (dashed line), 1.2 GeV (dashed-dotted line), and 1.3 GeV (double-dotted-dashed line).
@@
For the present numerical calculations, we have used the MMHT nucleonic parton distribution functions (PDFs) parameterization Harland-Lang:2014zoa in the MSbar scheme.
@@ III.1 Quasielastic scattering
Figure 10: $\sigma$ vs. $E_{\bar{\nu}_{\mu}}$ for the $\Lambda$ production. Solid (dashed) line represents the result using $M_{A}=1.026~{}(1.2)$ GeV. Experimental results: triangle right Eichten:1972bb , triangle up Erriquez:1977tr , square Erriquez:1978pg , triangle down ($\sigma=2.6^{+5.9}_{-2.1}\times 10^{-40}cm^{2}$anourakis:1980si , circle SKAT:1989nel are shown with error bars. Theoretical curves are of Kuzmin and Naumov Kuzmin:2008zz  (double dashed-dotted line), Brunner et al. SKAT:1989nel  (dashed line), Erriquez et al. Erriquez:1978pg  (dashed-double dotted line) obtained using Cabibbo theory with axial-vector dipole mass as 1 GeV, 1.1 GeV and 1 GeV, respectively, while the results of Wu et al. Wu:2013kla  (dotted line) and Finjord and Ravndal Finjord:1975zy  (dashed dotted line) are obtained using quark model.
@@
The numerical results for the differential and total cross sections are obtained using the BBBA05 Bradford:2006yz parameterization of the Sachs’ electric and magnetic form factors for the isovector vector form factors $f_{1,2}(Q^{2})$, axial vector form factor having dipole parameterization i.e. $g_{1}(Q^{2})=\frac{g_{1}(0)}{\left(1+\frac{Q^{2}}{M_{A}^{2}}\right)^{2}}$, $g_{1}(0)~{}(=1.26)$ is the axial charge and $M_{A}$ is the axial dipole mass. In the charged current sector, where PCAC is assumed to be valid, the pseudoscalar form factor $g_{3}(Q^{2})$, is dominated by the pion pole dominance of the divergence of the axial vector current (PDDAC) and is given in terms of $g_{1}(Q^{2})$ using the Goldberger-Trieman relation. However, $g_{3}(Q^{2})$ does not contribute in the limit $m_{l}\rightarrow 0$. As discussed above, the vector form factors are related to the electromagnetic form factors, which are determined precisely using the data from the real and virtual photon experiments on the nucleon targets, while the pseudoscalar form fIactor is related to the axial vector form factor. Thus, one of main source of uncertainty in the quasielastic scattering cross sections is associated with the uncertainty in the axial vector form factor. The value of axial vector charge $g_{1}(0)$ is determined precisely from the neutron decay experiments, and the value of $M_{A}$ is the major source of uncertainty as the various neutrino experiments report value of $M_{A}$ in the range 1–1.3 GeV (see Table 2.2 of Ref. SajjadAthar:2022pjt for details). Therefore, to show the dependence of $\frac{d\sigma}{dQ^{2}}$ on $M_{A}$, in Fig. 5, we have presented the results for the $Q^{2}$-distribution at $E_{{\nu_{\mu}}({\bar{\nu}_{\mu}})}=2$ GeV i.e. $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$, corresponding to the different values of $M_{A}$ i.e., $M_{A}=1.026$, 1.1, 1.2 and 1.3 GeV, for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced quasielastic scattering processes on the neutron ($\nu_{\mu}+n\rightarrow\mu^{-}+p$) and proton ($\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+n$) targets, respectively. It may be observed from the figure that the cross section increases with increase in the value of $M_{A}$, and this increase is more pronounced in the case of neutrino induced reaction than antineutrino induced reaction. An increment in the value of $M_{A}$ by 30% leads to an enhancement of about 25–35% in the cross section in the region of $Q^{2}\sim 0.4-0.6$ GeV2, for both neutrino and antineutrino induced reactions.
@@
In Figs. 6 and 7, we have presented the results for the $Q^{2}$-distribution at $E_{{\nu_{\mu}}({\bar{\nu}_{\mu}})}=2$ GeV for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced quasielastic scattering processes on the neutron ($\nu_{\mu}+n\rightarrow\mu^{-}+p$) and proton ($\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+n$) targets, respectively. The form of the induced tensor form factor $g_{2}(Q^{2})$, which is associated with the second class currents, is taken to be of dipole form with $g_{2}(0)=\pm 3$ and same $M_{A}$ as for the $g_{1}(Q^{2})$. The results presented in Fig. 6 are obtained to study the effect of second class currents assuming T-invariance by varying $g_{2}^{R}(0)$ between 00 and ±$\pm 3$, while the results have been presented in Fig. 7 to study the effect of second class currents assuming time reversal violation by varying $g_{2}^{I}(0)$ between 00 and ±$\pm 3$.
It may be observed from the figure that in the presence of second class current assuming time reversal invariance (Fig. 6), which requires $g_{2}(Q^{2})$ to be purely real i.e. $g_{2}(Q^{2})=g_{2}^{R}(Q^{2})$, the cross sections increases with the increase in the value of $g_{2}^{R}(0)$ and a change in the value of $g_{2}^{R}(0)$ from 0 to 3 results in an increase in the cross section by about $25\%$ at $Q^{2}\sim 0.4-0.6$ GeV2, for both neutrino and antineutrino induced reactions. We find that a change of sign in $g_{2}(0)$ i.e. $g_{2}(0)=-3$ gives almost the same result of the cross section as the one at $g_{2}(0)=+3$. In the case of second class currents assuming time reversal violation, which requires $g_{2}(Q^{2})$ to be purely imaginary i.e. $g_{2}(Q^{2})=g_{2}^{I}(Q^{2})$, a similar effect on the cross section (Fig. 7) is observed, as seen in the case of second class currents assuming time reversal invariance.
In Figs. 8 and  9, we have presented the results for the $Q^{2}$-distribution at $E_{\bar{\nu}_{\mu}}=2$ GeV for the charged current $\bar{\nu}_{\mu}$ induced $|\Delta S|=1$ quasielastic scattering processes on the proton ($\bar{\nu}_{\mu}+p\rightarrow\mu^{+}+\Lambda$) (left panel) and neutron ($\bar{\nu}_{\mu}+n\rightarrow\mu^{+}+\Sigma^{-}$) (right panel) targets, respectively. Assuming SU(3) symmetry, the $N-Y(=\Lambda,\Sigma)$ transition form factors have been taken to be of the similar form as taken in the case of the $|\Delta S|=0$ quasielastic scattering processes and discussed above. The results presented in Fig. 8 are obtained in the presence of the second class currents assuming time reversal invariance by varying $g_{2}^{R}(0)$ between 00 and ±$\pm 3$, while the results of T-violation studies have been presented in Fig. 9 by varying $g_{2}^{I}(0)$ between 00 and ±$\pm 3$. It may be observed from the figures that the non-zero value of the form factor associated with the second class currents assuming time reversal invariance as well as its violation, increases the cross section, which is more pronounced in the case of $\Lambda$ production as compared to the results obtained for $\Sigma^{-}$ production. Unlike the case of $\Delta S=0$ quasielastic processes, in the case of $\Lambda$ production, we find that the cross section obtained using the negative value of $g_{2}^{R}(0)$ is larger than the one obtained using the positive value of $g_{2}^{R}(0)$. Quantitatively, in the case of $\Lambda$ production, we find an enhancement of about $25\%$ ($40\%$) at $Q^{2}=0.5$ GeV2 when $g_{2}^{R}(0)=+3~{}(-3)$ is taken in the numerical calculations, and this enhancement increases with the increase in the value of $Q^{2}$. It may be pointed out that the $\Lambda$ production is more suitable for the SCC or T-violation studies as their effect on $\Lambda$ polarization can be studied by analyzing the angular distribution of pions produced in $\Lambda\rightarrow p\pi^{-}$ decays SajjadAthar:2022pjt . It has been observed (not shown here) Fatima:2018tzs that the presence of SCC with or with T invariance increases the total cross section. Moreover, a higher value of $M_{A}$ also leads to an increase in the cross section. Presently, in the neutrino physics community, there is an ongoing debate since the last decade to fix the value of $M_{A}$ as the results of the different experiments require different values of $M_{A}$ in the Monte Carlo to explain the data. In this light of this discussion, we have proposed that the presence of second class currents, if any, would help in fixing the value of $M_{A}$ Fatima:2018tzs . Furthermore, we have also studied the polarization of the produced hyperons Fatima:2018tzs and have shown that the transverse component of the hyperon polarization, which lies perpendicular to the reaction plane, would be nonzero only in the presence of the SCC with T violation. Since, the DUNE experiment would use LArTPC type of detector, where the three-dimensional image of the interaction vertex can be reconstructed, therefore, at DUNE the study of transverse component of the polarized hyperon is feasible. To conclude, if it becomes possible to put some limit on the transverse component of $\Lambda$ polarization in the future, then the information on the SCC can be obtained and using this information on the SCC form factor, it becomes possible to contraint the value of $M_{A}$ for the neutrino experiments.
[…]